# Slice-100K: A Multimodal Dataset for Extrusion-based 3D Printing

### v2

@@
Abstract G-code (Geometric code) or RS-274 is the most widely used computer numerical control (CNC) and 3D printing programming language. G-code provides machine instructions for the movement of the 3D printer, especially for the nozzle, stage, and extrusion of material for extrusion-based additive manufacturing. Currently there does not exist a large repository of curated CAD models along with their corresponding G-code files for additive manufacturing. To address this issue, we present Slice-100K, a first-of-its-kind dataset of over 100,000 G-code files, along with their tessellated CAD model, LVIS (Large Vocabulary Instance Segmentation) categories, geometric properties, and renderings. We build our dataset from triangulated meshes derived from Objaverse-XL and Thingi10K datasets. We demonstrate the utility of this dataset by finetuning GPT-2 on a subset of the dataset for G-code translation from a legacy G-code format (Sailfish) to a more modern, widely used format (Marlin). Slice-100K will be the first step in developing a multimodal foundation model for digital manufacturing.
@@ 1   Introduction
In recent years, the integration of digital design and computer-aided manufacturing processes has led to groundbreaking innovations in the manufacturing sector [1, 2]. One of the most transformative technologies at this intersection is additive manufacturing or 3D printing, which enables the physical manufacturing of digital assets [3, 4]. 3D printing surpasses the limitations of traditional manufacturing techniques by enabling the creation of parts with complex geometric shapes [5, 6]. A commonly used 3D printing method is extrusion-based additive manufacturing [7, 8], often based on Fused Deposition Modeling (FDM) for manufacturing plastic or polymer parts. In this method, bits of thermoplastic material are sequentially extruded from a heated nozzle, which has three degrees of freedom. The nozzle moves in a flat 2D plane and builds up the desired shape layer-by-layer.
@@
A typical 3D printing process begins with creating a 3D model of the part in a computer-aided design (CAD) program. This CAD model is then usually exported as a triangulated mesh file (for example, STL, PLY, or OBJ). The triangulated model is then “sliced” into multiple layers based on the resolution or the layer height of the 3D printer. Each layer is then converted into sequence of programmatic instructions for the movement of the 3D printer nozzle and extrusion of material along the boundary or “contour” of each layer. The instructions also include the movement of the nozzle and extrusion of material inside the contours or the “infill”. These instructions are then directly sent to the 3D printer for physical manufacturing. The most common representation for storing this information is G-code (Geometric code) or RS-274, a computer numerical control (CNC) programming language. G-code provides machine instructions for the movement of the 3D printer, especially for the nozzle, stage, and extrusion of material for extrusion-based additive manufacturing. Although some extensions of G-code have been written to include basic abstractions such as for-loops, the vast majority of G-code in use consists mainly of low-level instructions that provide a sequence of commands to be carried out by the 3D printer.
@@
Since 3D printing is a layered manufacturing process, it requires performing the slicing process. The slicing process operates on the entire object and splits it along the print direction (usually the Z-axis by default). Each layer is then used to generate the printer instructions for contour and infill. However, manual tuning of slicing software is usually required to achieve a good quality of the fabricated model. The iterative improvement of a given G-code file to produce a 3D-printed model that exactly matches its CAD representation is a non-trivial challenge. In addition, there are several “flavors” of G-code files depending on the compatibility of the 3D printer controller hardware. Due to the low-level nature of G-code, manually debugging a G-code file is cumbersome, if not impossible. Features such as line-level and layer-level natural language comments are very, very rare. While custom solutions such as regular expression matching could be leveraged for correcting G-code, they fall under a rigid set of methods and are not generalizable.
In the last few years, while advances in AI have impacted various domains, their potential in computer-aided design (CAD) and cybermanufacturing remains largely untapped. Modern LLMs and Vision-Language Models (VLMs) could provide an avenue to realize this potential. A LLMs’ ability to process, comprehend, and generate natural language descriptions, code, and other text data can be leveraged to interpret, generate, and manipulate G-code. LLMs for 3D shape modeling have been shown to enable operations on meshes [9, 10] and point clouds [11, 12]. G-code, with its unique language-based structure, presents distinct challenges for machine learning, mainly due to the context window limitations of current LLMs. Many existing deep-learning-based computer vision applications leverage 2D datasets (images), text descriptions, or a combination of such modalities for both supervised or self-supervised pre-training for foundation models. However, none of these datasets provide a curated avenue for training a manufacturing domain-specific foundation model.
To bridge this gap, we introduce Slice-100K, a curated multimodal dataset (see Figure 2 for reference) of G-code, CAD models, and renderings to facilitate the application of VLMs for additive manufacturing. We believe Slice-100K will encourage the research community to address new problems in the design and manufacturing space. Our dataset, built using models from Objaverse-XL and the Thingi10k dataset, encompasses a diverse range of 3D printable objects and provides a comprehensive resource for training a manufacturing domain-specific foundation model.
Contributions: This paper introduces Slice-100K, a multimodal dataset for manufacturing applications. The main features include a first-of-its-kind curated dataset of more than 100,000 G-code files along with their corresponding STL CAD files, renderings, LVIS (Large Vocabulary Instance Segmentation) categories, and metadata. We demonstrate the utility of Slice-100K by evaluating existing LLMs for G-code geometric transformations. We also showcase a novel application of our dataset on LLM-based G-code flavor translation. We believe that this multimodal dataset will be the starting point for a foundation model in digital manufacturing.
@@ 2   Background and Related Work
G-code: G-code forms a crucial intermediary between digital design and physical manufacturing, providing an expressive language-based representation for 3D objects. For example, the most straightforward command in G-code is G1 which directs the 3D printer to move its nozzle towards a spatial coordinate. This is usually followed by a coordinate in the form Xaaa Yaaa, where movement along the X and Y axes are given by a specific numeric value aaa. For extrusion-based 3D printers, a thermoplastic material is extruded from a heated nozzle that has three degrees of freedom. An example extrusion move is given by G1 X50.6 Y36.2 E2.3, where the nozzle moves 50.6 units along X, 36.2 units along Y and extrudes 2.3 units of material. Other commands instruct the printer to change settings such as the material/ink feed rate, or perform more complex movements without extruding material.
Language Models for Code: LLMs have been also used for programming language analysis and code generation. Coding-focused LLMs are mainly trained on a mix of web-scraped data, coding repositories, and instructions and often surpass general-purpose LLMs in code-related tasks. Current research has lead to many such models [13, 14, 15, 16, 17, 18]. Most notable ones include WizardCoder [18], Code Llama [19], and Instruct-CodeGen [16]. Codex [20] is an early model deployed under Github’s Copilot feature and acts as an IDE assistant that can understand local code context, make suggestions, and generate entire blocks of code.
3D Datasets: The current research community has proposed and leveraged various 3D datasets [21, 22, 23, 24, 25, 26, 27, 28]. Notable ones include Objaverse 1.0 [23] and Objaverse-XL [24], consisting of over 800K 3D models with higher quality textures and geometry types. The latter is a massive dataset of over 10 million objects gathered from various sources, including Thingi10K and GitHub repositories. The diversity of objects in terms of shapes and categories is an advantage for Objaverse-XL. Most of the datasets currently used by the research community provide a single modality (meshes or voxels), and some include text descriptions and renderings for visual supervision tasks. However, none of the currently available datasets provide curated assets for encouraging research in the manufacturing domain. The largest public G-code dataset we are aware of is the Greater G-code [29] dataset, which only contains 860 G-code files paired with their STL renderings.
@@
LLMs and 3D Datasets: Language understanding methods have been applied in the 3D domain for a wide array of tasks including 3D captioning [26, 30], object grounding [31, 11], 3D conversation [32], and text-conditioned generation [9, 10, 10]. Recently, there has been a surge of interest in multimodal large language models (MLLMs). MLLMs combine the language-based reasoning and knowledge of LLMs with the ability to comprehend other data modalities. Vision-augmented LLMs [33, 34, 35] encode images into an LLM’s embedding space. These methods have been subsequently extended to the 3D domain for different forms of 3D representation, such as pointclouds [12, 36], and sparse outdoor LiDAR data [37]. Paschalidou et al. [38] use a transformer-based model (not LLM) to autoregressively predict 3D objects in a scene. 3DLLM [11] maps 3D scenes to a set of 2D image embeddings and uses a query-token embedding technique based on BLIP-2’s Q-Former [34] to perform a diverse set of 3D-related tasks. GPT4Point [36] also leverages a similar Q-Former for point-text feature alignment. Chat3D [32] uses an object-centric 3D representation to train a 3D-LLM for dialogue. Feng et al. [39] does in-context learning on room layouts from the 3D-Front dataset [40]. PointBERT [41] did some early work on point-cloud representation learning with transformers. Fu et al. [30] align visual features from 3D scenes with text to finetune a LLaMa-2-chat-70B [42] model for scene understanding and question answering.
LLMs for Design and Manufacturing: Recent research has shown that natural language descriptions can be used for various tasks related to 3D printing, such as generating novel shapes [43, 44, 45, 46], editing scenes [47], and reasoning about geometry in the volume space [48]. Makatura et al. [49] thoroughly examine GPT-4’s suitability for automated design and manufacturing. Badini et al. [50] use ChatGPT to modify G-code, but they only alter the parameters in the G-code header. These modifications allow them to address common errors in the 3D printing process, such as warping, bed detachment, and stringing. Kulits et al. [51] train an LLM to autoregressively generate structured representations of simple 3D objects from the CLEVR dataset [52].
@@ 3.1   Data Collection Process
The Objaverse-XL dataset comprises of 3D objects gathered from Github, Thingiverse, Smithsonian Institution, Polycam, and Sketchfab. We gather data from the Thingiverse subset of Objaverse-XL. Thingiverse is one of the largest online platforms consisting of user-generated digital designs and is particularly focused on 3D printable files, encouraging community interaction and collaboration. A majority of these files are provided in the STL format and are available under Creative Commons licenses. The models on Thingiverse cover a wide range of categories, including functional parts, artistic creations, and educational tools. This extensive and diverse collection makes it an invaluable resource for creating comprehensive datasets for additive manufacturing.
The Thingi10K dataset [27] is a collection of 10,000 3D models sourced from Thingiverse. It is specifically curated for research purposes and provides a diverse set of models that are manifold and oriented, making them ideal for various computational geometry and 3D printing research applications. The dataset includes metadata and annotations that facilitate the development of machine learning models and other computational tools.
@@
G-code Generation: The G-code generation process is a critical component of the Slice-100K dataset. We utilize PrusaSlicer’s [53] command line functionality to slice all our models. Prusa’s slicer is an open-source and widely-used slicing software that prepares 3D models for printing by converting them into G-code, which provides specific instructions for 3D printers. Additionally, it allows for extensive configuration options, allowing for fine-tuning of print settings such as layer height, infill density, and support structures. This flexibility ensures that the generated G-code is high quality and suitable for different 3D printers and printing conditions. Furthermore, to minimize our data footprint, we generate G-code files in the binary G-code .bgcode format, a functionality recently incorporated by Prusa’s slicer. An important aspect of the slicing pipeline is the infill pattern selection, primarily due to its impact on total print time and structural properties of manufactured models. To encourage diversity among our G-code files with respect to structural properties, while slicing each STL file, we randomly select from four different infill patterns: (1) Gyroid: Empirically known to give equal strength across all directions and optimizes for a quicker print time; (2) Honeycomb: Uses a grid of hexagons, providing increased mechanical resistance, and non-crossing paths; (3) Cubic: Introduces crossing paths, potentially generating air-pockets; and (4) Grid: Uses a two-way checkerboard-like for faster infill.
@@
STL Renderings: To generate renderings of our STL files, we utilize Blender [54] rendering scripts made available by Objaverse-XL. We modify the script to generate a total of 10 views for each object - 6 orthogonal views (front, back, top, bottom, left, right) and 4 isometric views (captured from top four corner of a cube). Each object is rendered with a random color. These renderings are utilized for object category generation.
LVIS Object Categories: To generate the text category of each model in Slice-100K, we use the framework shown in Figure 3. For each model in our dataset, we assign the top 3 of the 1200+ LVIS (Large Vocabulary Instance Segmentation) categories [55]. This process helps enhance the utility of the dataset, enabling better categorization and facilitating more effective use in various research and development applications.
For each CAD model, we generate multiple views using Blender. This step ensures comprehensive visual coverage of the model, capturing its geometry from various angles. We input the renderings to a pre-trained Vision-Language model to generate image embeddings. We utilized the pre-trained CLIP-ViT-L-14 [56, 57] to create the image embeddings for each view. To integrate the information from multiple views, we computed the average embedding for each object. This average embedding combines the features from all views into a single, unified representation, providing a comprehensive summary of the visual characteristics of the object.
Parallel to the image embeddings, we also process the 1200+ LVIS categories to obtain the text embeddings for all categories. Using the average embedding, we then match each object to the closest categories in the text embedding. By comparing the average embeddings, we identify the top 3 most relevant LVIS categories for each object in our dataset.

### v3

@@
Abstract
G-code (Geometric code) or RS-274 is the most widely used computer numerical control (CNC) and 3D printing programming language. G-code provides machine instructions for the movement of the 3D printer, especially for the nozzle, stage, and extrusion of material for extrusion-based additive manufacturing. Currently, there does not exist a large repository of curated CAD models along with their corresponding G-code files for additive manufacturing. To address this issue, we present Slice-100K, a first-of-its-kind dataset of over 100,000 G-code files, along with their tessellated CAD model, LVIS (Large Vocabulary Instance Segmentation) categories, geometric properties, and renderings. We build our dataset from triangulated meshes derived from Objaverse-XL and Thingi10K datasets. We demonstrate the utility of this dataset by finetuning GPT-2 on a subset of the dataset for G-code translation from a legacy G-code format (Sailfish) to a more modern, widely used format (Marlin). Our dataset can be found here. Slice-100K will be the first step in developing a multimodal foundation model for digital manufacturing.
Figure 1: The Slice-100K dataset consists of STL files and their G-code counterparts. Each pair here consists of STL (left) and its slices (right) for G-code.
@@ 1 Introduction
In recent years, the integration of digital design and computer-aided manufacturing processes has led to groundbreaking innovations in the manufacturing sector (Alcácer and Cruz-Machado, 2019; Spanos et al., 2019). One of the most transformative technologies at this intersection is additive manufacturing or 3D printing, which enables the physical manufacturing of digital assets (Frank et al., 2019; Dilberoglu et al., 2017). 3D printing surpasses the limitations of traditional manufacturing techniques by enabling the creation of parts with complex geometric shapes (Pereira et al., 2019; Achillas et al., 2017). A commonly used 3D printing method is extrusion-based additive manufacturing (Gibson et al., 2021; Kumar, 2020), often based on Fused Deposition Modeling (FDM) for manufacturing plastic or polymer parts. In this method, bits of thermoplastic material are sequentially extruded from a heated nozzle, which has three degrees of freedom. The nozzle moves in a flat 2D plane and builds up the desired shape layer-by-layer.
@@
A typical 3D printing process begins with creating a 3D model of the part in a computer-aided design (CAD) program. This CAD model is then usually exported as a triangulated mesh file (for example, STL, PLY, or OBJ). The triangulated model is then “sliced” into multiple layers based on the resolution or the layer height of the 3D printer. Each layer is then converted into a sequence of programmatic instructions for the movement of the 3D printer’s nozzle and extrusion of material along the boundary or “contour” of each layer. The instructions also include the movement of the nozzle and extrusion of material inside the contours or the “infill.” These instructions are then directly sent to the 3D printer for physical manufacturing. The most common representation for storing this information is G-code (Geometric code) or RS-274, a computer numerical control (CNC) programming language. G-code provides machine instructions for the movement of the 3D printer, especially for the nozzle, stage, and extrusion of material for extrusion-based additive manufacturing. Although some extensions of G-code have been written to include basic abstractions such as for-loops, the vast majority of G-code in use consists mainly of low-level instructions that provide a sequence of commands to be carried out by the 3D printer.
Since 3D printing is a layered manufacturing process, it requires performing the slicing process. The slicing process operates on the entire object and splits it along the print direction (usually the Z-axis by default). Each layer is then used to generate the printer instructions for contour and infill. However, achieving high-quality fabricated models often requires manual tuning of the slicing software. The iterative improvement of a given G-code file to produce a 3D-printed model that exactly matches its CAD representation is a non-trivial challenge. In addition, there are several “flavors” of G-code files depending on the compatibility of the 3D printer’s controller hardware. Due to the low-level nature of G-code, manually debugging a G-code file is cumbersome, if not impossible. Features such as line-level and layer-level natural language comments are infrequent. While custom solutions such as regular expression matching could be leveraged for correcting G-code, they fall under a rigid set of methods and are not generalizable.
In the last few years, while advances in AI have impacted various domains, their potential in computer-aided design (CAD) and cybermanufacturing remains largely untapped. Modern LLMs and Vision-Language Models (VLMs) could provide an avenue to realize this potential. The ability of LLMs to process, comprehend, and generate natural language descriptions, code, and other text data can be leveraged to interpret, generate, and manipulate G-code. LLMs for 3D shape modeling have been shown to enable operations on meshes (Siddiqui et al., 2023; Yin et al., 2023) and point clouds (Hong et al., 2023; Xu et al., 2023). G-code, with its unique language-based structure, presents distinct challenges for machine learning, mainly due to the context window limitations of current LLMs. Many existing deep-learning-based computer vision applications leverage 2D datasets (images), text descriptions, or a combination of such modalities for both supervised or self-supervised pre-training of foundation models (see Table 1). However, none of these datasets provide a curated avenue for training a manufacturing domain-specific foundation model.
To bridge this gap, we introduce Slice-100K, a curated multimodal dataset (see Figure 2 for reference) of G-code, CAD models, renderings, and geometric properties to facilitate the application of VLMs for additive manufacturing. We believe Slice-100K will encourage the research community to address new problems in the design and manufacturing space. Our dataset, built using models from Objaverse-XL and the Thingi10k dataset, encompasses a diverse range of 3D printable objects and provides a comprehensive resource for training a manufacturing domain-specific foundation model.
Contributions: This paper introduces Slice-100K, a multimodal dataset for manufacturing applications. The main features include a first-of-its-kind curated dataset of more than 100,000 G-code files along with their corresponding STL CAD files, renderings, LVIS (Large Vocabulary Instance Segmentation) categories, and geometric properties. We demonstrate the utility of Slice-100K by evaluating existing LLMs for G-code geometric transformations. We also showcase a novel application of our dataset on LLM-based G-code flavor translation. We believe that this multimodal dataset will be the starting point for a foundation model in digital manufacturing.
@@ 2 Background and Related Work
G-code: G-code forms a crucial intermediary between digital design and physical manufacturing, providing an expressive language-based representation for 3D objects. For example, the most straightforward G-code command is G1, which directs the 3D printer to move its nozzle towards a spatial coordinate. This is usually followed by a coordinate in the form Xaaa Yaaa, where movement along the X and Y axes are given by a specific numeric value aaa. For extrusion-based 3D printers, a thermoplastic material is extruded from a heated nozzle that has three degrees of freedom. An example extrusion move is given by G1 X50.6 Y36.2 E2.3, where the nozzle moves 50.6 units along X, 36.2 units along Y and extrudes 2.3 units of material. Other commands instruct the printer to change settings, such as the material/ink feed rate, or perform more complex movements without extruding material.
Language Models for Code: LLMs have also been used for programming language analysis and code generation. Coding-focused LLMs are mainly trained on a mix of web-scraped data, coding repositories, and instructions and often surpass general-purpose LLMs in code-related tasks. Current research has lead to many such models (Allal et al., 2023; Christopoulou et al., 2022; Shen et al., 2023; Chaudhary, 2024; Achiam et al., 2023; Luo et al., 2023a). Most notable ones include WizardCoder (Luo et al., 2023a), Code Llama (Roziere et al., 2023), and Instruct-CodeGen (Chaudhary, 2024). Codex (Chen et al., 2021) is an early model deployed under Github’s Copilot feature and acts as an IDE assistant that can understand local code context, make suggestions, and generate entire blocks of code.
3D Datasets: The current research community has proposed and leveraged various 3D datasets (Collins et al., 2022; Chang et al., 2015; Deitke et al., 2023a, b; Koch et al., 2019; Luo et al., 2023b; Zhou and Jacobson, 2016; Uy et al., 2019). Notable ones include Objaverse 1.0 (Deitke et al., 2023a) and Objaverse-XL (Deitke et al., 2023b), with the former consisting of over 800K 3D models with higher quality textures and geometry types (Table 1). The latter is a massive dataset of over 10 million objects gathered from various sources, including Thingi10K and GitHub repositories. The diversity of objects in terms of shapes and categories is an advantage for Objaverse-XL. Most of the datasets currently used by the research community provide a single modality (meshes or voxels), and some include text descriptions and renderings for visual supervision tasks. However, none of the currently available datasets provide curated assets for encouraging research in the manufacturing domain. The largest public G-code dataset we are aware of is the Greater G-code (Issak, 2022) dataset, which only contains 860 G-code files paired with their STL renderings.
@@
LLMs and 3D Datasets: Language understanding methods have been applied in the 3D domain for a wide array of tasks including 3D captioning (Luo et al., 2023b; Fu et al., 2024), object grounding (Achlioptas et al., 2020; Hong et al., 2023), 3D conversation (Wang et al., 2023), and text-conditioned generation (Siddiqui et al., 2023; Yin et al., 2023, 2023). Recently, there has been a surge of interest in multimodal large language models (MLLMs). MLLMs combine the language-based reasoning and knowledge of LLMs with the ability to comprehend other data modalities. Vision-augmented LLMs (Liu et al., 2023; Li et al., 2023a; Zhu et al., 2023) encode images into an LLM’s embedding space. These methods have been subsequently extended to the 3D domain for different forms of 3D representation, such as point clouds (Xu et al., 2023; Qi et al., 2024), and sparse outdoor LiDAR data (Yang et al., 2023). Paschalidou et al. (2021) use a transformer-based model (not LLM) to predict 3D objects in a scene autoregressively. 3DLLM (Hong et al., 2023) maps 3D scenes to a set of 2D image embeddings and uses a query-token embedding technique based on BLIP-2’s Q-Former (Li et al., 2023a) to perform a diverse set of 3D-related tasks. GPT4Point (Qi et al., 2024) also leverages a similar Q-Former for point-text feature alignment. Chat3D (Wang et al., 2023) uses an object-centric 3D representation to train a 3D-LLM for dialogue. Feng et al. (2023) does in-context learning on room layouts from the 3D-FRONT dataset (Fu et al., 2021). PointBERT (Yu et al., 2022) did some early work on point-cloud representation learning with transformers. Fu et al. (2024) align visual features from 3D scenes with text to finetune a LLaMa-2-chat-70B (Touvron et al., 2023) model for scene understanding and question answering.
LLMs for Design and Manufacturing: Recent research has shown that natural language descriptions can be used for various tasks related to 3D printing, such as generating novel shapes (Sanghi et al., 2022; Marshall et al., 2023; Jain et al., 2022; Lin et al., 2023), editing scenes (Haque et al., 2023), and reasoning about geometry in the volume space (Kerr et al., 2023). Makatura et al. (2023) thoroughly examine GPT-4’s suitability for automated design and manufacturing. Badini et al. (2023) use ChatGPT to modify G-code, but they only alter the parameters in the G-code header. These modifications allow them to address common errors in the 3D printing process, such as warping, bed detachment, and stringing. Kulits et al. (2024) train an LLM to autoregressively generate structured representations of simple 3D objects from the CLEVR dataset (Johnson et al., 2017).
@@ 3.1 Data Collection Process
The Objaverse-XL dataset comprises of 3D objects gathered from Github, Thingiverse, Smithsonian Institution, Polycam, and Sketchfab. We collect data from the Thingiverse subset of Objaverse-XL. Thingiverse is one of the largest online platforms consisting of user-generated digital designs and is particularly focused on 3D printable files, encouraging community interaction and collaboration. A majority of these files are provided in the STL format and are available under Creative Commons licenses. The models on Thingiverse cover a wide range of categories, including functional parts, artistic creations, and educational tools. This extensive and diverse collection makes it an invaluable resource for creating comprehensive datasets for additive manufacturing.
The Thingi10K dataset (Zhou and Jacobson, 2016) is a collection of 10,000 3D models sourced from Thingiverse. It is curated explicitly for research purposes and provides a diverse set of models that are manifold and oriented, making them ideal for various computational geometry and 3D printing research applications. The dataset includes metadata and annotations that facilitate the development of machine learning models and other computational tools.
@@
G-code Generation: The G-code generation process is a critical component of the Slice-100K dataset. We utilize PrusaSlicer’s (PrusaSlicer, 2023) command line functionality to slice all our models. Each model is sliced using two distinct G-code flavors—Sailfish and Marlin. Prusa’s slicer is an open-source and widely-used slicing software that prepares 3D models for printing by converting them into G-code, which provides specific instructions for 3D printers. Additionally, it allows for extensive configuration options, allowing for fine-tuning of print settings such as layer height, infill density, and support structures. This flexibility ensures that the generated G-code is high quality and suitable for different 3D printers and printing conditions. Furthermore, to minimize our data footprint, we generate G-code files in the binary G-code (.bgcode) format, a functionality recently incorporated by Prusa’s slicer. An important aspect of the slicing pipeline is the infill pattern selection, primarily due to its impact on total print time and structural properties of manufactured models. To encourage diversity among our G-code files with respect to structural properties, while slicing each STL file, we randomly select from four different infill patterns: (1) Gyroid: Empirically known to give equal strength across all directions and optimizes for a quicker print time; (2) Honeycomb: Uses a grid of hexagons, providing increased mechanical resistance, and non-crossing paths; (3) Cubic: Introduces crossing paths, potentially generating air-pockets; and (4) Grid: Uses a two-way checkerboard-like pattern for faster infill.
@@
STL Renderings: To generate renderings of our STL files, we utilized Blender (Org, 2024) rendering scripts made available by Objaverse-XL. We modified the scripts to generate a total of 10 views for each object - 6 orthogonal views (front, back, top, bottom, left, right) and 4 isometric views (captured from top four corners of a cube). Each object was rendered with a random color. These renderings were utilized for object category generation.
LVIS Object Categories: To generate the text category of each model in Slice-100K, we use the framework shown in Figure 3. For each model in our dataset, we assign the top 3 of the 1200+ LVIS (Large Vocabulary Instance Segmentation) categories (Gupta et al., 2019). This process helps enhance the utility of the dataset, enabling better categorization and facilitating more effective use in various research and development applications. For each CAD model, we generate multiple views using Blender. This step ensures comprehensive visual coverage of the model, capturing its geometry from various angles. The generated renderings serve as input to a pre-trained Vision-Language model to generate image embeddings. Specifically, we utilize a pre-trained CLIP-ViT-L-14 (Radford et al., 2021; OpenAI, 2024) to obtain embeddings for each view. To integrate information from multiple views, we compute an average embedding for each object. This average embedding combines the features from all views into a single, unified representation, providing a comprehensive summary of the visual characteristics of the object.
In parallel, we also process the 1200+ LVIS categories to obtain the text embeddings for all categories. Using the average embedding, we then match each object to the closest categories in the text embedding. By comparing the average embeddings, we identify the top 3 most relevant LVIS categories for each object in our dataset.